OpenStudy (anonymous):

Solve. 8 = 2^(x + 4)

OpenStudy (across):

\[8=2^{x+4}\]\[2^3=2^{x+4}\]\[3=x+4\]\[x=-1\]

OpenStudy (experimentx):

2^3=2^(x+4) 3=x+4 x=-1